OpenStudy (anonymous):

Stoichiometry help please... 2Al+3H2SO4 -> Al2(SO4)3+ 3H2 If you have 78.3 grams of Al how many liters of H2 will be produced?

OpenStudy (anonymous):

\[Moles:Al=78.3/27=2.9mol\]Therefore:\[Moles:H _{2}=2.9\times \frac{3}{2}=4.35mol\]Assuming 1 mole of any gas occupies 24L at room temperature and pressure, the volume of hydrogen produced is:\[4.35\times24=104.4L\]